OpenStudy (anonymous):

Tess and Dan have $24.00 each to spend at a book fair, where all students receive a 15% discount. They both want to purchase a copy of the same book, which normally sells for $24.50 plus 10% sales tax. To check if she has enough to purchase the book, Tess takes 15% of $24.50 and subtracts that amount from the normal price. She takes 10% of the discounted selling price and adds it back to find the purchase amount. Dan takes 85% of the normal purchase price and then computes 110% of the reduced price. Is Tess correct? Is Dan correct? Do they have enough money to purchase the book? Explain your a

OpenStudy (anonymous):

Plz Help

OpenStudy (texaschic101):

Tess ... 15% of 24.50 = 0.15(24.5) = 3.675 -- this is your 15% discount subtracts from normal price = 24.50 - 3.675 = 20.825 -- this is reduced price 10% of discounted selling price = .10(20.825) = 2.0825 adds back to find purchase price = 20.825 + 2.0825 = 22.9075 rounds to 22.91 ==================== Dan ... 85% of normal purchase price = 0.85(24.5) = 20.825 110% of reduced price = 1.10 (20.825) = 22.9075 rounds to 22.91 ==================== Is Tess correct ? Yes Is Dan correct ? Yes Do they have enough money ? Yes Basically, subtracting 15% is the same as paying 85%...and finding 10% of selling price and adding that back to discount price to find purchase price is the same as adding 110% of reduced price to get purchase price.